# Verifiable Fine-Tuning for LLMs: Zero-Knowledge Training Proofs Bound to Data Provenance and Policy

### v1

@@ V-B Verification Performance
Per-step prover time on Tier-P2 with r=8r{=}8, 2,048 tokens averages $16.8\pm 0.9$ s; for r=16r{=}16 it is $31.2\pm 1.4$ s. Peak memory is $23.6\pm 1.1$ GB and $37.9\pm 1.5$ GB, respectively. Constraint counts scale near-linearly with $P_{\mathrm{PEFT}}$ and superlinearly with sequence length due to attention checks; the slope is lower than naive quadratic because our lookup-based softmax imposes row-sum constraints rather than polynomial expansions per entry. Fig. 4 (referenced) shows a breakdown: witness generation consumes $\sim$35% of time, polynomial/FFT evaluation $\sim$28%, MSM/FRI $\sim$30%, and overheads the remainder. Compared to training-proof prototypes that circuitize dense updates [9, 11], our PEFT restriction reduces constraint volume by an order of magnitude for typical transformer blocks.
@@ V-C Policy and Privacy Outcomes
We estimate index leakage by training an adversary to predict whether a given example index appeared in a hidden batch using only proof-level signals and public constants. The AUROC is $0.500\pm 0.004$, consistent with chance, and mutual information estimates are within numerical noise. Padding step-proof time to fixed windows eliminates residual timing side channels. These findings align with the zero-knowledge claims of our backend and improve upon non-ZK audit trails, which may leak through metadata [35].
@@
All runs pass constant-equality checks against $h_{\Pi}$; we injected controlled deviations (learning-rate spikes, altered $\beta_{2}$) and observed immediate proof failure, confirming that optimizer semantics and schedule constraints are enforced. This concrete guarantee complements high-level provenance capture [4, 10]: while provenance systems record intended configurations, VFT proves the executed trajectory adhered to them.
@@ V-D Ablations
Sweeping the softmax error budget $\delta_{\mathrm{sm}}$ from $1.5{\times}10^{-3}$ to $6.0{\times}10^{-3}$ reduces prover time by $10.8\%$ but increases PPL by $0.7$ absolute (Mix-A). Fixed-point fractional bits $q\in\{12,16,20\}$ show the expected trade-off: moving from q=14q{=}14 to q=12q{=}12 saves $7.2\%$ prover time but increases ECE by $0.8\%$; moving to q=20q{=}20 yields negligible utility gains but raises constraint counts $\sim$12%. Table III summarizes. These trends mirror broader ZK-ML observations that the right approximation frontier is problem-specific [14].
@@
Increasing LoRA rank from r=4r{=}4 to r=16r{=}16 improves instruction metrics by $\sim$0.6 points but doubles prover cost; placing adapters in attention and MLP layers yields modest quality gains (+0.2+0.20.30.3 points) for $\sim$25% higher cost. A normalized metric—proof-time per trainable parameter—is nearly constant across ranks, indicating linearity in $P_{\mathrm{PEFT}}$. This supports our design choice of PEFT-restricted circuits and resonates with recent practice that parameter-efficient fine-tuning dominates real deployments [44, 43, 47].
@@
Moving from 2,048 to 4,096 tokens increases constraint counts by $2.23\times$ and prover time by $2.38\times$ at fixed rank; PPL improves slightly due to longer contexts. Attention matmuls drive the increase; lookup-based softmax keeps the growth sub-quadratic relative to naive polynomialization. Fig. 7 (referenced) shows log–log slopes and confidence intervals. These costs outpace those reported for inference-only ZK [15] because training proofs include backward passes and optimizer updates.
@@
Private-sampling windows add $6.4\%$ prover time for membership proofs and quota attestations, and increase leaf proof size by $\sim$18 KB. In return, they provide index privacy without impeding verifiability. Interleaving public windows enables third-party spot checks; the mix has negligible effect on quality. This hybrid design extends PoL-informed thinking [1, 19] by binding the sampler to both reproducible and privacy-preserving regimes.
@@
With K=200K{=}200 clients, verifying $q{=}10\%$ of clients’ steps per round detects injected anomalies (learning-rate tampering, out-of-policy sampling) with $99\%$ probability within two rounds; $q{=}5\%$ detects within three rounds. Aggregator-side epoch proofs include a coverage certificate; bandwidth overhead remains acceptable because per-step proofs are optional and folded when provided. These findings echo proposals for probabilistic audits in decentralized LLM training [20] and demonstrate composability with succinct proofs.
@@ V-E Comparisons to Prior Art
Classical PoL requires trajectory evidence but often relies on logging schemes that can be cherry-picked or manipulated [1]. Incentive-secure variants push further but still stop short of numerical verification [19]. VFT subsumes PoL’s spirit while providing strong substance: hash-chained states (Fig. 1) tied to step-wise ZK proofs of updates. In practice, VFT’s per-epoch proofs and public/private sampling windows retain the operational advantages of PoL—committing early, verifying late—while adding cryptographic guarantees.
@@
Compared to inference-focused systems [3, 15], VFT addresses the training gap by proving optimizer transitions and per-batch membership. Against proof-of-training prototypes [9, 11], we observe better accuracy–cost trade-offs at scale because (i) PEFT confines proof obligations and (ii) approximation budgets are tuned to language modeling numerics, following zkLLM lessons [13]. Surveys emphasize recursion and error accounting as key enablers [6, 8]; our results empirically validate both.
@@
Provenance capture and automated screenings are necessary but not sufficient [4, 35]. VFT binds provenance claims to computation-level evidence: the manifest’s quotas and constants become circuit-enforced invariants, and violations are cryptographically impossible to hide. Distributional attestation and licensing audits benefit directly: regulators can inspect bin-level counters without accessing private examples [7, 2].
@@
Watermarks are complementary post-release signals [26, 33]. However, their brittleness under adaptive attacks [28, 29, 32] and mixed success in practice [27, 34] make them insufficient for compliance. Our manifest can record where (if at all) watermarks are inserted, aiding downstream detectors; but the training story rests on proofs, not beacons. Robust watermark designs [30, 31] help on the output side and can be paired with VFT for defense-in-depth.
@@
Decentralized and edge learning has seen sustained progress on communication and personalization [37, 50, 43, 47, 46, 44, 52]. VFT complements these lines with verifiability: the aggregator can accept only updates accompanied by partial proofs or coverage certificates; clients can hide indices in private-sampling mode. Our experiments demonstrate feasibility within the bandwidth and latency envelopes typical of enterprise deployments.
@@ V-F Visual Analyses and Qualitative Insights
Fig. 6 visualizes per-epoch bin utilization under $\gamma$. Heatmaps reveal that telemetry-tagged bins stay well below ceilings by design; safety bins operate near 70–80% utilization, balancing exposure and overfitting risks. This visualization style, borrowed from data-governance dashboards [4], translates naturally into audit artifacts: inspectors can accept or challenge quota choices with clear context.
@@ V-G Case Studies
We simulate a policy that restricts medical-tagged items to at most 5% per epoch. VFT enforces and proves compliance; a post-hoc audit requests selective disclosure of aggregate token counts in medical bins. The auditor verifies proofs in <200<200 ms and accepts the run without accessing raw examples. This workflow operationalizes licensing/attribution recommendations from large-scale audits [2] and extends distributional attestation methods with end-to-end training claims [7].
@@
In a K=200K{=}200 setting, one client injects out-of-policy telemetry descriptors. The aggregator’s probabilistic audits [20] flag the mismatch within a round; subsequent aggregation excludes tainted updates until the client attests a corrected manifest. Because client batches are private-sampled, no sensitive indices are revealed during remediation; yet the zero-knowledge attestations suffice to restore trust.
@@ V-H Failure Modes and Diagnostics
Setting $\delta_{\mathrm{sm}}$ too tight causes prover slowdowns and memory spikes; too loose raises PPL and ECE. Table III includes a red zone where constraints explode ($\delta_{\mathrm{sm}}\leq 5{\times}10^{-4}$) or fidelity suffers ($\delta_{\mathrm{sm}}\geq 8{\times}10^{-3}$). The sweet spot aligns with zkLLM design guidance [13].
@@ V-I Security Reflection
Our threat model enumerates attempts to alter hyperparameters, skip steps, or exceed quotas. Injections of learning-rate spikes or modified $\beta_{2}$ in AdamW are consistently rejected by the circuit constraints. Attempts to sneak in out-of-policy examples fail at the membership/quota checks. Efforts to exploit approximation tolerances (e.g., crafting logits near clipping thresholds) do not succeed because bounds are enforced per-step and hash-chained; such attacks would need to violate either the commitment binding or proof soundness. These guarantees are stronger than watermark-based accountability [26, 28] and complementary to provenance logs [35].
@@ V-J Ethical Considerations in Results
Because VFT can enforce quotas on sensitive bins, it could be misused to exclude protected domains. Our evaluation emphasizes transparent manifests that make such policies explicit and auditable. In federated settings, clients maintain autonomy over private sampling; auditors can request aggregate counters without coercive disclosure. This design balances accountability and privacy, resonating with principles advocated by large-scale data audits [2].
@@ V-L Summary of Findings
VFT delivers cryptographically strong, privacy-preserving assurances about the fine-tuning trajectory while preserving model utility within tight budgets. Compared to the state of the art in proof-of-training and zk-LLM [9, 11, 13], our PEFT-centric circuits and approximation budgeting reduce prover cost and maintain accuracy. By binding provenance manifests to computation-level proofs, VFT operationalizes calls for accountable data governance [2, 4, 35] and sidesteps the brittleness of watermark-only accountability [26, 34]. Federated evaluations confirm compatibility with heterogeneous devices and probabilistic audits [37, 43, 20]. Across datasets and settings, quota compliance is perfect, index leakage is negligible, and verification costs fit within the latency envelopes of practical audits. These results support the view that verifiable fine-tuning can be deployed today in centralized and decentralized pipelines to close the trust gap around LLM adaptation.

### v2

@@ V-B Verification Performance
Per-step prover time on Tier-P2 with r=8r{=}8, 2,048 tokens averages $16.8\pm 0.9$ s; for r=16r{=}16 it is $31.2\pm 1.4$ s. Peak memory is $23.6\pm 1.1$ GB and $37.9\pm 1.5$ GB, respectively. Constraint counts scale near-linearly with $P_{\mathrm{PEFT}}$ and superlinearly with sequence length due to attention checks; the slope is lower than naive quadratic because our lookup-based softmax imposes row-sum constraints rather than polynomial expansions per entry. Fig. 4 (referenced) shows a breakdown: witness generation consumes $\sim$35% of time, polynomial/FFT evaluation $\sim$28%, MSM/FRI $\sim$30%, and overheads the remainder. Compared to training-proof prototypes that circuitize dense updates [sun2023zkdl, abbaszadeh2024zkpot], our PEFT restriction reduces constraint volume by an order of magnitude for typical transformer blocks.
@@ V-C Policy and Privacy Outcomes
We estimate index leakage by training an adversary to predict whether a given example index appeared in a hidden batch using only proof-level signals and public constants. The AUROC is $0.500\pm 0.004$, consistent with chance, and mutual information estimates are within numerical noise. Padding step-proof time to fixed windows eliminates residual timing side channels. These findings align with the zero-knowledge claims of our backend and improve upon non-ZK audit trails, which may leak through metadata [schlegel2025isProvenance].
@@
All runs pass constant-equality checks against $h_{\Pi}$; we injected controlled deviations (learning-rate spikes, altered $\beta_{2}$) and observed immediate proof failure, confirming that optimizer semantics and schedule constraints are enforced. This concrete guarantee complements high-level provenance capture [schelter2024provenanceScreen, schlegel2023mlflow2prov]: while provenance systems record intended configurations, VFT proves the executed trajectory adhered to them.
@@ V-D Ablations
Sweeping the softmax error budget $\delta_{\mathrm{sm}}$ from $1.5{\times}10^{-3}$ to $6.0{\times}10^{-3}$ reduces prover time by $10.8\%$ but increases PPL by $0.7$ absolute (Mix-A). Fixed-point fractional bits $q\in\{12,16,20\}$ show the expected trade-off: moving from q=14q{=}14 to q=12q{=}12 saves $7.2\%$ prover time but increases ECE by $0.8\%$; moving to q=20q{=}20 yields negligible utility gains but raises constraint counts $\sim$12%. Table III summarizes. These trends mirror broader ZK-ML observations that the right approximation frontier is problem-specific [taylorZKML2025].
@@
Increasing LoRA rank from r=4r{=}4 to r=16r{=}16 improves instruction metrics by $\sim$0.6 points but doubles prover cost; placing adapters in attention and MLP layers yields modest quality gains (+0.2+0.20.30.3 points) for $\sim$25% higher cost. A normalized metric—proof-time per trainable parameter—is nearly constant across ranks, indicating linearity in $P_{\mathrm{PEFT}}$. This supports our design choice of PEFT-restricted circuits and resonates with recent practice that parameter-efficient fine-tuning dominates real deployments [10251628, yao2024ferrari, yan2024peaches].
@@
Moving from 2,048 to 4,096 tokens increases constraint counts by $2.23\times$ and prover time by $2.38\times$ at fixed rank; PPL improves slightly due to longer contexts. Attention matmuls drive the increase; lookup-based softmax keeps the growth sub-quadratic relative to naive polynomialization. Fig. 7 (referenced) shows log–log slopes and confidence intervals. These costs outpace those reported for inference-only ZK [wang2023ezdps] because training proofs include backward passes and optimizer updates.
@@
Private-sampling windows add $6.4\%$ prover time for membership proofs and quota attestations, and increase leaf proof size by $\sim$18 KB. In return, they provide index privacy without impeding verifiability. Interleaving public windows enables third-party spot checks; the mix has negligible effect on quality. This hybrid design extends PoL-informed thinking [jia2021proof, zhao2024polIncentive] by binding the sampler to both reproducible and privacy-preserving regimes.
@@
With K=200K{=}200 clients, verifying $q{=}10\%$ of clients’ steps per round detects injected anomalies (learning-rate tampering, out-of-policy sampling) with $99\%$ probability within two rounds; $q{=}5\%$ detects within three rounds. Aggregator-side epoch proofs include a coverage certificate; bandwidth overhead remains acceptable because per-step proofs are optional and folded when provided. These findings echo proposals for probabilistic audits in decentralized LLM training [probAudit2025] and demonstrate composability with succinct proofs.
@@ V-E Comparisons to Prior Art
Classical PoL requires trajectory evidence but often relies on logging schemes that can be cherry-picked or manipulated [jia2021proof]. Incentive-secure variants push further but still stop short of numerical verification [zhao2024polIncentive]. VFT subsumes PoL’s spirit while providing strong substance: hash-chained states (Fig. 1) tied to step-wise ZK proofs of updates. In practice, VFT’s per-epoch proofs and public/private sampling windows retain the operational advantages of PoL—committing early, verifying late—while adding cryptographic guarantees.
@@
Compared to inference-focused systems [liu2021zkcnn, wang2023ezdps], VFT addresses the training gap by proving optimizer transitions and per-batch membership. Against proof-of-training prototypes [sun2023zkdl, abbaszadeh2024zkpot], we observe better accuracy–cost trade-offs at scale because (i) PEFT confines proof obligations and (ii) approximation budgets are tuned to language modeling numerics, following zkLLM lessons [sun2024zkllm]. Surveys emphasize recursion and error accounting as key enablers [zkvml2023survey, peng2025zkmlsurvey]; our results empirically validate both.
@@
Provenance capture and automated screenings are necessary but not sufficient [schelter2024provenanceScreen, schlegel2025isProvenance]. VFT binds provenance claims to computation-level evidence: the manifest’s quotas and constants become circuit-enforced invariants, and violations are cryptographically impossible to hide. Distributional attestation and licensing audits benefit directly: regulators can inspect bin-level counters without accessing private examples [dold2024attesting, longpre2024data].
@@
Watermarks are complementary post-release signals [kirchenbauer2023watermark, wen2023treering]. However, their brittleness under adaptive attacks [gu2024learnability, jovanovic2024wmstealing, zhang2024mipSteal] and mixed success in practice [dathathri2024natureWatermark, liu2024textwmSurvey] make them insufficient for compliance. Our manifest can record where (if at all) watermarks are inserted, aiding downstream detectors; but the training story rests on proofs, not beacons. Robust watermark designs [zhao2023provable, qu2025multibit] help on the output side and can be paired with VFT for defense-in-depth.
@@
Decentralized and edge learning has seen sustained progress on communication and personalization [liu2021communication, liu2023yoga, yao2024ferrari, yan2024peaches, 10508334, 10251628, liu2024enhancing]. VFT complements these lines with verifiability: the aggregator can accept only updates accompanied by partial proofs or coverage certificates; clients can hide indices in private-sampling mode. Our experiments demonstrate feasibility within the bandwidth and latency envelopes typical of enterprise deployments.
@@ V-F Visual Analyses and Qualitative Insights
Fig. 6 visualizes per-epoch bin utilization under $\gamma$. Heatmaps reveal that telemetry-tagged bins stay well below ceilings by design; safety bins operate near 70–80% utilization, balancing exposure and overfitting risks. This visualization style, borrowed from data-governance dashboards [schelter2024provenanceScreen], translates naturally into audit artifacts: inspectors can accept or challenge quota choices with clear context.
@@ V-G Case Studies
We simulate a policy that restricts medical-tagged items to at most 5% per epoch. VFT enforces and proves compliance; a post-hoc audit requests selective disclosure of aggregate token counts in medical bins. The auditor verifies proofs in <200<200 ms and accepts the run without accessing raw examples. This workflow operationalizes licensing/attribution recommendations from large-scale audits [longpre2024data] and extends distributional attestation methods with end-to-end training claims [dold2024attesting].
@@
In a K=200K{=}200 setting, one client injects out-of-policy telemetry descriptors. The aggregator’s probabilistic audits [probAudit2025] flag the mismatch within a round; subsequent aggregation excludes tainted updates until the client attests a corrected manifest. Because client batches are private-sampled, no sensitive indices are revealed during remediation; yet the zero-knowledge attestations suffice to restore trust.
@@ V-H Failure Modes and Diagnostics
Setting $\delta_{\mathrm{sm}}$ too tight causes prover slowdowns and memory spikes; too loose raises PPL and ECE. Table III includes a red zone where constraints explode ($\delta_{\mathrm{sm}}\leq 5{\times}10^{-4}$) or fidelity suffers ($\delta_{\mathrm{sm}}\geq 8{\times}10^{-3}$). The sweet spot aligns with zkLLM design guidance [sun2024zkllm].
@@ V-I Security Reflection
Our threat model enumerates attempts to alter hyperparameters, skip steps, or exceed quotas. Injections of learning-rate spikes or modified $\beta_{2}$ in AdamW are consistently rejected by the circuit constraints. Attempts to sneak in out-of-policy examples fail at the membership/quota checks. Efforts to exploit approximation tolerances (e.g., crafting logits near clipping thresholds) do not succeed because bounds are enforced per-step and hash-chained; such attacks would need to violate either the commitment binding or proof soundness. These guarantees are stronger than watermark-based accountability [kirchenbauer2023watermark, gu2024learnability] and complementary to provenance logs [schlegel2025isProvenance].
@@ V-J Ethical Considerations in Results
Because VFT can enforce quotas on sensitive bins, it could be misused to exclude protected domains. Our evaluation emphasizes transparent manifests that make such policies explicit and auditable. In federated settings, clients maintain autonomy over private sampling; auditors can request aggregate counters without coercive disclosure. This design balances accountability and privacy, resonating with principles advocated by large-scale data audits [longpre2024data].
@@ V-L Summary of Findings
VFT delivers cryptographically strong, privacy-preserving assurances about the fine-tuning trajectory while preserving model utility within tight budgets. Compared to the state of the art in proof-of-training and zk-LLM [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm], our PEFT-centric circuits and approximation budgeting reduce prover cost and maintain accuracy. By binding provenance manifests to computation-level proofs, VFT operationalizes calls for accountable data governance [longpre2024data, schelter2024provenanceScreen, schlegel2025isProvenance] and sidesteps the brittleness of watermark-only accountability [kirchenbauer2023watermark, liu2024textwmSurvey]. Federated evaluations confirm compatibility with heterogeneous devices and probabilistic audits [liu2021communication, yao2024ferrari, probAudit2025]. Across datasets and settings, quota compliance is perfect, index leakage is negligible, and verification costs fit within the latency envelopes of practical audits. These results support the view that verifiable fine-tuning can be deployed today in centralized and decentralized pipelines to close the trust gap around LLM adaptation.